# The Illusion of Readiness: Stress Testing Large Frontier Models on Multimodal Medical Benchmarks

### v2

@@ IV. Toward Community Standards for Evaluation Reform
GPT-5 83.33 51.67 -31.66
Gemini-2.5 Pro 80.83 47.50 -33.33
OpenAI-o3 76.67 52.50 -24.17
o4-mini 71.67 37.50 -34.17
GPT-4o 36.67 41.67 +5.00
Note: Models with genuine visual understanding should revise their prediction when the image is changed to support a different answer. Instead, performance drops sharply—especially for top-performing models—suggesting reliance on shortcut visual–answer associations rather than robust image interpretation.
(a) Reasoning complexity rubric, scored along five sub-axes reflecting depth of inference, uncertainty, and clinical decision impact.
(b) Visual complexity rubric, capturing task reliance on detailed image interpretation, spatial understanding, and visual comparison.
Figure 6: Clinician-defined rubrics for evaluating benchmark complexity. These rubrics enable systematic profiling of benchmarks along clinically relevant reasoning and visual dimensions, revealing variation that standard metadata overlooks.
@@
Note: We report quantitative results for GPT-4V, MAIRA-2, and MedVERSA based on the ReXrank leaderboard [20], which provides standardized evaluation for AI-powered chest X-ray radiology report generation.
Appendix A Robustness Score Computation
@@
Appendix B More stress test results
@@
Figures 7-9 show representative examples of stress test 2 and 3 (Section 2). These cases demonstrate how models often rely on textual shortcuts: when visual information is removed, most models still produce seemingly correct answers by exploiting superficial textual cues rather than genuine multimodal reasoning. In contrast, GPT-4o is the only model that consistently abstains or selects “Unknown” in the absence of the image.
@@
Figures  10-12 illustrate stress test 5, which probe visual comprehension under visual substitution. A model with true visual understanding should revise its prediction to align with the substituted image. Instead, performance collapses across all models, revealing that models’ reliance on shortcut strategies and their limited ability to incorporate visual evidence, whereas accurate performance would require aligning predictions with the substituted image. Once shortcut strategies are disrupted, their true comprehension is far weaker than the benchmark scores suggest.
@@
Note: Based on the modality and anatomical location of the original problem image, we retrieved candidate distractor images using Google Image Search [21]. These candidates were visually similar to the original but depicted different conditions. To ensure validity, each image was manually reviewed to verify consistency between the visual content and the associated disease label.
Figure 7: Illustration of model behavior under Stress Tests 2 and 3: Case 1. We apply three modifications to the original image-based multiple-choice question (ground truth: Dermatomyositis): (i) remove the image, (ii) shuffle the answer options, and (iii) replace one option with “Unknown.” When the image is absent, the appropriate behavior is to abstain (or to select “Unknown” if available). However, most models still provide a diagnosis—often repeating Dermatomyositis even after the options are shuffled—whereas GPT-4o is the only model that consistently refuses to answer without the image.
Figure 8: Illustration of model behavior under Stress Tests 2 and 3: Case 2. We apply three modifications to the original image-based multiple-choice question (ground truth: Chiari I malformation): (i) remove the image, (ii) shuffle the answer options, and (iii) replace one option with “Unknown.” When the image is absent, the expected behavior is to abstain (or to select “Unknown” if available). Nevertheless, most models continue to provide a diagnosis—often repeating the original ground truth—whereas GPT-4o is the only model that consistently either abstains or selects “Unknown” in the absence of the image.
Figure 9: Illustration of model behavior under Stress Tests 2 and 3: Case 3. We apply three modifications to the original image-based multiple-choice question (ground truth: Pneumatosis cystoides coli): (i) remove the image, (ii) shuffle the answer options, and (iii) replace one option with “Unknown.” When the image is absent, the expected behavior is to abstain (or to select “Unknown” if available). Nevertheless, most models continue to provide a diagnosis—often repeating the original ground truth—whereas GPT-4o is the only model that consistently refuses to answer without the image.
Appendix C More reasoning failure results
@@
As discussed in Section 2, we identify several key patterns of reasoning failure: incorrect logic despite correct answers, reinforced visual misunderstandings, and uninformative reasoning steps. In this appendix, we present additional illustrative cases of these failure modes in Figures 1315. These cases further demonstrate how such reasoning errors manifest across different tasks and inputs.
Appendix D Evaluation on report generation task
@@
Extented Data Table 7 summarizes the report generation performance on the MIMIC-CXR test set for both general-purpose LMMs and specialized medical models (e.g., MAIRA-2 [22], MedVERSA [23]). Evaluation is conducted with five complementary metrics: (1) $1/\text{RadCliQ-v1}$, where higher values indicate better clinical correctness per the RadCliQ composite metric [24]; (2) BLEU Score, measuring n-gram overlap with reference reports [25]; (3) BERT Score, assessing semantic similarity via contextual embeddings [26]; (4) Semantic Score, capturing higher-level alignment of generated and reference report content [27]; and (5) RadGraph F1, evaluating extraction accuracy of clinical entities and relations from generated reports [28]. Across all metrics, MedVERSA achieves the highest performance, particularly in semantic alignment and clinical information extraction. Its advantage stems from domain-specific training on paired image–text data, which enables mastery of radiological terminology, fine-grained visual patterns, and clinical reasoning. In contrast, general-purpose LMMs lack such specialization, limiting their ability to capture subtle radiographic findings.
@@
Among the general-purpose models, GPT-4o and Gemini-2.5 Pro achieve the strongest performance, approaching the task-specific MAIRA-2; nevertheless, both still lag substantially behind MedVERSA, which maintains a clear lead across all metrics. Notably, GPT-5 and o3—despite substantially outperforming GPT-4o on VQA-style benchmarks—do not exhibit corresponding gains in report generation. This divergence underscores that improvements in general multimodal VQA do not necessarily translate into better radiology report generation, a task that demands precise identification of clinical findings, faithful spatial grounding, and disciplined use of domain-specific terminology.
Appendix E Prompt Design
@@
This appendix details the prompt formulations used in Sec 2, which were designed to evaluate model performance across different tasks. The report generation prompt follows the structured medical reporting paradigm described in recent work [29].
@@
Please only answer with the index and content of the option.

### v3

@@ Discussion
GPT-5 83.33 [75.83,90.00] 51.67 [42.50,60.83] -31.66
Gemini-2.5 Pro 80.83 [73.33,87.50] 47.50 [38.33,56.67] -33.33
OpenAI-o3 76.67 [69.17,84.17] 52.50 [43.33,61.67] -24.17
OpenAI-o4-mini 71.67 [63.33,79.17] 37.50 [29.17,45.83] -34.17
GPT-4o 36.67 [28.33,45.00] 41.67 [33.33,50.83] +5.00
Note: Models with genuine visual understanding should revise their prediction when the image is changed to support a different answer. Instead, performance drops sharply, especially for top-performing models, suggesting reliance on shortcut visual–answer associations rather than robust image interpretation.
Table 8: Reasoning with CoT Accuracy (%) of models on filtered subsets of NEJM (118 items) and VQA-RAD (100 items) under Without CoT and With CoT conditions. $\Delta$ denotes the change in accuracy (With CoT - Without CoT).
Model NEJM VQA-RAD
W/o CoT W/ CoT $\Delta$ W/o CoT W/ CoT $\Delta$
GPT-5 81.36 [73.73,88.14] 79.66 [72.03,86.44] -1.70 65.00 [55.00,74.00] 66.00 [57.00,75.00] +1.00
Gemini-2.5 Pro 83.90 [77.12,89.83] 81.36 [73.73,88.14] -2.54 67.00 [58.00,76.00] 64.00 [55.00,73.00] -3.00
OpenAI-o3 81.36 [73.73,88.14] 79.66 [72.03,86.44] -1.70 66.00 [57.00,75.00] 63.00 [54.00,72.00] -3.00
OpenAI-o4-mini 73.73 [65.25,81.36] 79.66 [72.03,86.44] +5.93 59.00 [49.00,69.00] 61.00 [51.00,71.00] +2.00
GPT-4o 67.80 [59.32,76.27] 65.25 [55.93,73.73] -2.55 63.00 [54.00,72.00] 60.00 [50.00,70.00] -3.00
Notes: Two-decimal precision preserved from raw logs; $\Delta$ computed as exact subtraction (two decimals).
Table 9: Reasoning complexity rubric. Clinician-defined rubric scored along five sub-axes capturing depth of inference, uncertainty handling and clinical decision impact.
Criteria Description Rubric levels Score
Reasoning steps How many thinking steps are needed? One-step (e.g., identify organ)
Combine findings
Differential diagnosis or staged inference
1
2
3
Clinical context Does the task require clinical background or patient context? None
Some (e.g., complaint)
Full scenario (history, labs, etc.)
1
2
3
Time / sequence information Does the task depend on change over time (e.g., prior scans, progression)? No
Sometimes
Always or essential
1
2
3
Uncertainty handling Does the task involve ambiguity? No
Sometimes vague
Requires hedging or considering multiple possibilities
1
2
3
Decision impact Does the outcome affect clinical decisions? No impact (descriptive / trivia)
Some impact (e.g., triage, follow-up)
High impact (diagnosis or treatment guiding)
1
2
3
Table 10: Visual complexity rubric. Clinician-defined rubric capturing reliance on detailed image interpretation, spatial understanding and cross-view or temporal comparison.
Criteria Description Rubric levels Score
Text-alone solvable Could the task be answered without the image? Yes
Maybe
No , image essential
1
2
3
Visual detail needed How carefully must the image be examined? Glance is enough
Moderate inspection
Expert-level detail (e.g., subtle lesion, boundary)
1
2
3
Spatial understanding Is spatial localization important? Not needed
Side or region matters
Exact location critical
1
2
3
Image-quality sensitivity Would small variations in the image alter the answer? No
Somewhat
Yes , subtle differences matter
1
2
3
Multi-view / temporal comparison Does the task require comparing across views or timepoints? No comparison needed
Sometimes referenced
Critical comparison across views or timepoints
1
2
3
@@
Note: We report quantitative results for GPT-4V, MAIRA-2, and MedVERSA from ReXrank leaderboard, which provides standardized evaluation for AI-powered chest X-ray radiology report generation.
Table 12: Abstention rate(%) of models on JAMA (1,314 items), NEJM (947 items) and NEJMstress (175 items) under Image+Text and Text Only conditions.
Model JAMA NEJM NEJMstress
Img+Txt Txt Only Img+Txt Txt Only Img+Txt Txt Only
GPT-5 0 0 0 0 0 0
Gemini-2.5 Pro 0 0 0 0 0 0
OpenAI-o3 13.17 0 21.54 0 0 0
OpenAI-o4-mini 12.02 0 11.19 0 0 0
GPT-4o 0 0 0 45.2 0 91.43
Table 13: Structured taxonomy of failure modes in multimodal medical reasoning.
Domain Failure Mode Cognitive Mechanism Clinical Consequence Revealed by
Input Handling Perceptual hallucination / omission Extrinsic hallucination (misreads or invents visual findings) False positives (unnecessary tests) or false negatives (missed lesions) T2, T5
Refusal miscalibration Uncertainty misalignment (answers when unsafe or abstains when needed) Unsafe guessing or withholding guidance T2–T4
Modality omission Faithfulness violation (ignores available input modality) Missed findings; incomplete assessment T1–T2
Reasoning and Inference Shortcut reliance Heuristic bias / gaming (pattern use over reasoning) Unstable performance under distribution shift T3–T5
Fabricated reasoning Extrinsic hallucination (unsupported or false explanation) Misleads clinician; false interpretability T2, T6
Inconsistency Logical incoherence (reasoning conflicts with evidence or answer) Confusion; reduced auditability of outputs T1–T6
Output Communication Misleading fluency Fluency illusion (language masks factual errors) Hard-to-detect errors; over-trust in output T2, T6
Unsafe recommendation / omission Information hazard (harmful or incomplete suggestion) Potential diagnostic or treatment error T5, T6
Supplementary Note 1 | Robustness Score Computation
@@
Supplementary Note 2 | More stress test results
@@
Supplementary Figs. 1–3 show representative examples of stress test 2 and 3. These cases demonstrate how models often rely on textual shortcuts: when visual information is removed, most models still produce seemingly correct answers by exploiting superficial textual cues rather than genuine multimodal reasoning. In contrast, GPT-4o is the only model that consistently abstains or selects “Unknown” in the absence of the image.
@@
Supplementary Figs. 4–6 illustrate stress test 5, which probe visual comprehension under visual substitution. A model with true visual understanding should revise its prediction to align with the substituted image. Instead, performance collapses across all models, revealing that models’ reliance on shortcut strategies and their limited ability to incorporate visual evidence, whereas accurate performance would require aligning predictions with the substituted image. Once shortcut strategies are disrupted, their true comprehension is far weaker than the benchmark scores suggest.
@@
Note: Based on the modality and anatomical location of the original problem image, we retrieved candidate distractor images using Google Image Search. These candidates were visually similar to the original but depicted different conditions. To ensure validity, each image was manually reviewed to verify consistency between the visual content and the associated disease label.
Supplementary Note 3 | More reasoning failure results
@@
We identify several key patterns of reasoning failure: incorrect logic despite correct answers, reinforced visual misunderstandings, and uninformative reasoning steps. In this appendix, we present additional illustrative cases of these failure modes in Supplementary Figs. 8-10. These cases further demonstrate how such reasoning errors manifest across different tasks and inputs.
Supplementary Note 4 | Evaluation on report generation task
@@
Supplementary Table 10 summarizes the report generation performance on the MIMIC-CXR test set for both general-purpose LMMs and specialized medical models (e.g., MAIRA-2, MedVERSA). Evaluation is conducted with five complementary metrics: (1) $1/\text{RadCliQ-v1}$, where higher values indicate better clinical correctness per the RadCliQ composite metric; (2) BLEU Score, measuring n-gram overlap with reference reports; (3) BERT Score, assessing semantic similarity via contextual embeddings; (4) Semantic Score, capturing higher-level alignment of generated and reference report content; and (5) RadGraph F1, evaluating extraction accuracy of clinical entities and relations from generated reports. Across all metrics, MedVERSA achieves the highest performance, particularly in semantic alignment and clinical information extraction. Its advantage stems from domain-specific training on paired image–text data, which enables mastery of radiological terminology, fine-grained visual patterns, and clinical reasoning. In contrast, general-purpose LMMs lack such specialization, limiting their ability to capture subtle radiographic findings.
@@
Among the general-purpose models, GPT-4o and Gemini-2.5 Pro achieve the strongest performance, approaching the task-specific MAIRA-2; nevertheless, both still lag substantially behind MedVERSA, which maintains a clear lead across all metrics. Notably, GPT-5 and OpenAI-o3, despite substantially outperforming GPT-4o on VQA-style benchmarks, do not exhibit corresponding gains in report generation. This divergence underscores that improvements in general multimodal VQA do not necessarily translate into better radiology report generation, a task that demands precise identification of clinical findings, faithful spatial grounding, and disciplined use of domain-specific terminology.
Supplementary Note 5 | Prompt Design
@@
This appendix details the prompt formulations used in stress-tests, which were designed to evaluate model performance across different tasks. The report generation prompt follows the structured medical reporting paradigm described in recent work.
@@
Provide answer with the index and content of the option, and place it within <answer></answer>
@@
Supplementary Note 6 | Model identifiers and versions
Table 14: Model identifiers and version information. Exact identifiers and release versions for all evaluated models, accessed through official APIs or public checkpoints as of August 2025.
Model Name Model Identifier Version / Release Date
GPT-4 gpt-4 turbo-2024-04-09
GPT-4o gpt-4o 2024-08-06
OpenAI-o3 OpenAI-o3 2025-04-16
OpenAI-o4-Mini OpenAI-o4-mini 2025-04-16
GPT-5 gpt-5 2025-08-07
DeepSeek-VL2 deepseek-ai/deepseek-vl2 API model (no version)
Gemini 2.5 Pro Preview (06-05) gemini-2.5-pro-preview-06-05 2025-06-05
Gemini 2.5 Pro Preview (03-25) gemini-2.5-pro-preview-03-25 2025-03-25